# Investigating and Mitigating Stereotype-aware Unfairness in LLM-based Recommendations

### v1

@@ 2.2. Evaluation of Stereotype-aware Fairness
Table 2. Comparison of pre-trained LLM and LLM-RS, where LLM-RS are LoRA fine-tuned on the MovieLens dataset.
@@
im. ex. cf. AUC Precis. Recall
LLM 61.43 62.83 53.71 62.23 62.64 81.67
LLM-RS 77.05 80.94 74.99 70.27 68.32 86.33
2.2.2. Stereotype-aware Fairness.
Following the above design of stereotype measurement, we can unify users and items into stereotype groups (i.e., two-sided groups), to address the simultaneous existence of LLM-encoded stereotypes in both user’s and item’s word embeddings. Different from previous fairness that individually considers user groups or item groups, we propose a new variant of fairness between two-sided groups, where each group contains a stereotype-oriented subset of both users and items.
In pursuit of fair recommendations against stereotypes, LLM-RS should not over-recommend items in any specific stereotype group $G$ compared to the proportion of G in user-item interactions. In particular, given a user in any specific stereotype group $G$, the recommendation proportion of target items in the same stereotype group $G$ should be calibrated to the proportion of $\mathbf{G}$ (i.e., $h_{u\in G}$) in the user’s interaction history.
Formally, we propose stereotype-aware fairness at the group level for each stereotype group $G\in\mathcal{G}$. Given the set of $N$ recommendations $\mathcal{S}=\{u_{i},\hat{v}_{i}\}_{i=1}^{N}$ where each tuple denotes the target item $\hat{v}$ (or $\hat{v}_{i}$) recommended to a user $u$ (or $u_{i}$), the evaluation metric of stereotype-aware fairness is defined as:
(4) $\bm{SF}=1-\frac{1}{|\mathcal{G}|}\sum_{G\in\mathcal{G}}{\frac{\sum_{u\in\mathcal{S}}{h_{u\in G}}}{\sum_{\hat{v}\in\mathcal{S}}{\mathbb{1}(d_{\hat{v}\in G})}}},$
where $\bm{SF}>0$ implies that LLM-RS amplify the recommendation proportion of any stereotype group $G$ that a user belongs to (i.e., $u\in G$) compared to the proportion of $G$ in user-item interactions, as marked by over-recommendation in Figure 1. Similarly, $\bm{SF}<0$ indicates under-recommendation.
2.3. Analysis of Stereotype-aware Fairness
2.3.1. Preliminary Experimental Setup
To validate the existence of the proposed stereotype-aware fairness in LLM-RS, exploratory experiments are conducted on two widely-used recommendation datasets: MovieLens and BookCrossing. The detailed description of datasets and evaluation metrics can be found in Section 4.1.
2.3.2. Analysis of RQ1
As illustrated in Table 1, we compare the performance and fairness of LLM-RS when performing recommendations between users and items from different stereotype groups. As for performance comparisons, the recommendation quality between users and items from inconsistent stereotype groups is significantly inferior to that of consistent stereotype groups. In particular, the AUC, Precision, and Recall of LLM-RS decrease by 30%, 19%, and 5%, respectively, in recommendations between users $u\in G$ and target items $v\notin G$, such as recommending romantic movies (e.g., female stereotype) to male users. These observations imply that the recommendation quality of LLM-RS is sensitive to stereotypes, emphasizing the concern of stereotype-aware fairness toward the trustworthiness of LLM-RS.
Despite the downgrade in performance, the fairness of LLM-RS regarding recommendations between inconsistent stereotype groups indicates a notably 67%-99% smaller value of $\bm{SF}$. In other words, the recommendation proportion of items from a stereotype group $G$ is much more calibrated to the proportion of $G$ in user-item interactions. However, negative $\bm{SF}$ can be observed in the MovieLens dataset, meaning that LLM-RS rarely recommend items $v\in G$ to users $u\notin G$ despite a large proportion of $G$ in the user’s historical interactions. The aforementioned differences in recommendations between consistent and inconsistent groups imply that LLM-RS tend to amplify the discrimination between user-side stereotypes and item-side stereotypes, such as exhibiting over-recommendation between users and items from the same stereotype group.
2.3.3. Analysis of RQ2
Since stereotypes are simultaneously embedded in the word embeddings of users and items (e.g., user profiles and item titles), we aim to investigate the degree of stereotype-aware fairness under different levels of stereotypes, namely implicit, explicit, and counterfactual settings (Xu et al., 2023). Specifically, the implicit setting only provides item titles in the input prompt of LLM-RS to infer the user-side stereotype without the actual user profile. In explicit and counterfactual settings, both user profiles and item titles are provided, as detailed in Section 4.1.4. By comparing results between implicit and explicit settings, the unfairness of LLM-RS with implicit stereotypes decreases by 18%-66% for inconsistent stereotype groups and 0.8%-9% for consistent stereotype groups. These results support our findings that stereotypes exist in the word embeddings of both users and items, leading to stereotype-aware fairness in LLM-RS. Notably, LLM-RS shows a reduction in unfairness by a percentage of 4%-7% in a counterfactual world of explicit stereotypes, highlighting the potential of utilizing different stereotypes (e.g., counterfactual stereotype) to develop effective methods in addressing stereotype-aware fairness in LLM-RS.
2.3.4. Ablation on LLM-encoded Stereotype w/wo Fine-tuning.
@@ 2.3.3. Analysis of RQ2
LLMs trained on large-scale uncurated data inherit stereotypes that are embedded ubiquitously in word embeddings. This leads to LLM-RS exhibiting stereotypical linguistic associations between users and items (e.g., user profiles and item titles). By fine-tuning LLMs on recommendation data (i.e., LLM-RS), we aim to investigate the influence of recommendation data to LLM-encoded stereotypes. As illustrated in Table 2, the performance of LLM-RS exceed LLM by 12.9%, 9%, and 5.7% in terms of AUC, Precision, and Recall, respectively, indicating the effectiveness of fine-tuning LLMs on recommendation data. As for the influence to stereotypes measured by stereotype-aware fairness, a significant downgrade of fairness can be observed, varying from 25% to 39% under different fairness settings. This implies that the stereotypes in word embeddings of LLM-RS can be amplified by fine-tuning on recommendation data, emphasizing the concern of stereotype-aware fairness toward the trustworthiness of LLM-RS.
Figure 3. The overall framework of the proposed MoS. In (a), multi-stereotype prompting elicits biases with respect to different stereotype groups. In (b), MoS mitigates the elicited stereotypes in recommendation tasks, where unbiased representations are generated and integrated with LLM-RS via soft prompts. In (c), effective learning objectives are designed to facilitate both the recommendation performance and the stereotype-aware fairness.
3. THE PROPOSED METHOD
@@
3.1. Overview of Mixture-of-Stereotypes (MoS)
As shown in Figure 3, the proposed framework consists of two key modules, namely multi-stereotype prompting and stereotype-wise routing, along with effective learning objectives. The first module aims to distinguish different stereotypes in the textual information of recommendation tasks. Due to the ubiquitous stereotypes embedded in the word embeddings of LLM-RS, it is necessary to effectively identify different stereotype groups in addressing stereotype-aware fairness. Accordingly, we introduce a multi-stereotype prompting module to elicit each stereotype group via textual prompts, as recent studies have revealed that stereotypes in LLMs can be modified by prompts (Guo et al., 2022; Yang et al., 2023; Li et al., 2024). Thereafter, the elicited stereotypes are captured by utilizing different stereotype-relevant experts (Cai et al., 2024) and encoded into soft prompts of LLM-RS via prompt tuning. Second, a stereotype-wise routing module is developed to learn unbiased representations that are consistent with the distribution of stereotypes in user-item interactions, utilizing a reweighting strategy over multiple stereotype-relevant experts and carefully designed learning objectives.
3.2. Multi-stereotype Prompting
As illustrated in Figure 3, a multi-stereotype prompting module is applied to the textual input prompt of recommendation tasks. The idea of multi-stereotype prompting is to amplify the LLM-encoded stereotypes in recommendation tasks, such as gender, age, and race. In particular, we design stereotype-biased prompts by inserting textual descriptions with respect to each stereotype group, as recent studies have revealed that stereotypes in LLMs can be modified by prompts (Guo et al., 2022; Yang et al., 2023; Li et al., 2024). For example, a prompting template for gender stereotypes is demonstrated as follows:
[Instruction] Given a female user’s interaction history, identify whether this female user will like the target item.
Formally, given a recommendation task prompt $x^{\mathrm{rec}}$ and a set of stereotype-biased prompts $\{x_{i}^{\mathrm{stereotype}}\}_{i=1}^{|\mathcal{G}|}$ (e.g., ”female”), the tokenization output of multi-stereotype prompting is formalized as
(5) $x^{\mathrm{rec}},\{x_{i}^{\mathrm{stereotype}}\}_{i=1}^{|\mathcal{G}|}\rightarrow\{(c^{\mathrm{rec}},c_{i}^{\mathrm{stereotype}})\}_{i=1}^{|\mathcal{G}|},$
@@ 3.2. Multi-stereotype Prompting
3.3. Stereotype-wise Routing
@@
(7) $p_{n}(\bm{c}_{i})=\begin{cases}[softmax(\{p_{k}(\bm{c}_{i})\}_{k\in\mathcal{K}})]_{n},&\text{if}\ n=k\\ 0,&\text{otherwise}\end{cases},$
where $k\in\mathcal{K}$ denotes the index of each activated expert in the top-K set $\mathcal{K}$ (i.e., $|\mathcal{K}|\leq N$). It is worth noting that the routing probabilities of inactivated experts are set to zero, since a reweighting strategy of the routing probabilities will be designed, as illustrated in Section 3.3.2. Intuitively, the zeroing operation reinforces the polarization of routing different stereotype information to different experts, facilitating the goal of stereotype-wise routing.
3.3.2. Reweighting Layer.
In light of our preliminary findings that the discrimination between user-side and item-side stereotypes can be alleviated in a counterfactual world, we explore combining multiple stereotypes (i.e., stereotype-relevant experts) to address stereotype-aware fairness in LLM-RS. In particular, we aim to generate unbiased representations that are consistent with the distribution of stereotypes in user-item interactions, by learning adaptive weights across different stereotype-relevant experts. As a natural solution, the learning objectives of adaptive weights can be obtained by stereotype-aware fairness, as illustrated in Eq. (4), and added to the training loss of LLM-RS. However, unlike discriminative recommendation models, it is challenging to update generative LLM-RS with a learning objective given by group-level fairness (Dai et al., 2024b; Atwood et al., 2024; Tommasel, 2024). To be specific, the learning objectives of generative LLM-RS are to maximize the likelihood of the label tokens of each output (e.g., target item), which intrinsically fall short in calculating an auxiliary loss over a group of outputs.
@@
(8) $\{w_{n}\}_{n=1}^{N}=\bm{R}(\frac{1}{|\mathcal{G}|}\sum_{i=1}^{|\mathcal{G}|}{\{p_{n}(\bm{c}_{i})\}_{n=1}^{N}}).$
In other words, $\{w_{n}\}_{n=1}^{N}$ implies a weighted average of multiple stereotypes encoded in different stereotype-relevant experts.
3.3.3. Prompt-based Experts.
With the aforementioned reweighting strategy, the mixture of stereotype-relevant experts can generate unbiased representations against stereotypes, taking advantage of the weighted average of different stereotypes that accord with the distribution of stereotypes in user-item interactions. To adapt the learned unbiased representations to LLM-RS, we design prompt-based experts to encode the generated representations as soft prompts, which can be seamlessly integrated with LLM-RS via prompt tuning (Lester et al., 2021). Formally, the soft prompts generated by experts $\{\bm{E}_{n}\}_{n=1}^{N}$ can be formalized as:
(9) $e=\sum_{n=1}^{N}{w_{n}\bm{E}_{n}(c^{\mathrm{rec}})}.$
@@ 3.4. Learning Objectives
(11) $\mathcal{L}_{\mathrm{rec}}=-\log\Pr_{\Phi+\Theta}(\hat{y}|e,c^{\mathrm{rec}}),$
@@
(12) $\mathcal{L}_{\mathrm{fair}}\coloneq\mathop{\min}_{\Theta}\|\bm{SF}\|.$
Notably, we further introduce a stereotype diversity regulation term to enhance the learning of stereotype information via different expert networks. In detail, expert parameter redundancy is invertible in multiple-expert architectures, leading to similar representations (i.e., learnable knowledge) across multiple expert networks (Liu et al., 2023; Tian et al., 2024; Dai et al., 2024a). Therefore, a stereotype diversity regulation term is designed to maximize the distance between the weights of each stereotype-relevant expert (i.e., diversified representations) as follows:
(13) $\mathop{\max}_{\{w_{n}\}_{n=1}^{N}}\{\mathcal{L}_{\mathrm{div}}\coloneq\mathop{\min}_{i\neq j}(\|w_{i}-w_{j}\|^{2})\}.$
@@
(14) $\mathcal{L}=\mathcal{L}_{\mathrm{rec}}+\mathcal{L}_{\mathrm{fair}}+\mathcal{L}_{\mathrm{div}}.$
4. EXPERIMENT
@@
4.1. Experiment Setup
4.1.1. Datasets
@@
MovieLens100K (Harper and Konstan, 2015) is a movie recommendation dataset, which provides user-movie interactions and textual information including movie titles and user profiles. In particular, we utilize the binary gender feature of users to assess the gender stereotype of LLM-RS.
BookCrossing (Ziegler et al., 2005) is a book recommendation dataset, which provides user-book interactions and textual information including book titles and user profiles. In particular, we utilize the age feature and divide users into teen and adult groups (i.e., under/beyond 18), to assess the age stereotype of LLM-RS.
To maintain a manageable dataset size for efficient LLM-RS training, similar to recent studies (Bao et al., 2023; Liao et al., 2023), we process the original datasets by randomly sampling 10,000 sequences (i.e., each contains 11 chronologically interactions). To construct sequential recommendation scenarios, we adopt the leave-one-out strategy and retain the first 10 items in each sequence as the historical interaction, and the last item as the target item. For both datasets, we split the data points of user-item interactions into training, validation, and testing sets with a ratio of 8:1:1, which prevents data leakage. The detailed statistics of experimental datasets are provided in Table 3.
Table 3. Basic statistics of experimental datasets.
@@
  • * $G_{1}$/$G_{2}$ denote the stereotype groups of male/female and adult/teen, respectively, in MovieLens and BookCrossing datasets.
4.1.2. Evaluation Metrics
We assess the proposed stereotype-aware fairness and the recommendation performance of LLM-RS.
Fairness. The proposed stereotype-aware fairness $\bm{SF}$ can be calculated according to Eq. (4). A smaller value of $\bm{SF}$ suggests a minor degree of over-recommendation between users and items from the same stereotype at the group level.
Table 4. Results of fairness-oriented methods. We report the average results over three independent runs with random seeds. The reported results are multiplied by 100, where boldface and underline indicate the best and second best score, respectively. The improvements of our proposed method are compared to the best baseline.

### v2

@@ V-C Analysis of RQ2
TABLE II: Comparison of pre-trained LLM and LLM-RS fine-tuned on the MovieLens dataset. The reported results are multiplied by 100, where boldface indicates the best score.
@@
implicit explicit counterfactual AUC Precision Recall
LLM 60.43 66.87 57.54 60.96 65.56 60.47
LLM-RS 75.22 79.17 72.09 67.16 75.83 69.62
V-D Ablation on LLM-encoded Stereotype w/wo Fine-tuning.
@@
LLMs trained on large-scale uncurated data inherit stereotypes that are embedded ubiquitously in word embeddings. This leads to LLM-RS exhibiting stereotypical linguistic associations between users and items (e.g., user profiles and item titles). By fine-tuning LLMs on recommendation data (i.e., LLM-RS), we aim to investigate the influence of recommendation data to LLM-encoded stereotypes. As illustrated in Table II, the performance of LLM-RS exceeds LLM by 11%, 15%, and 15% in terms of AUC, Precision, and Recall, respectively, indicating the effectiveness of fine-tuning LLMs on recommendation data. As for the influence on stereotypes measured by stereotype-aware fairness, a significant downgrade of fairness can be observed, varying from 25% to 39% under different fairness settings. This implies that the stereotypes in word embeddings of LLM-RS can be amplified by fine-tuning on recommendation data, emphasizing the concern of stereotype-aware fairness toward the trustworthiness of LLM-RS.
VI THE PROPOSED METHOD
@@
Figure 4: The overall framework of the proposed MoS. In (a), multi-stereotype prompting elicits biases with respect to different stereotype groups. In (b), MoS mitigates the elicited stereotypes in recommendation tasks, where unbiased representations are generated and integrated with LLM-RS via soft prompts. In (c), effective learning objectives are designed to facilitate both the recommendation performance and the stereotype-aware fairness.
VI-A Overview of Mixture-of-Stereotypes (MoS)
As shown in Figure 4, our framework consists of three key modules, namely multi-stereotype prompting and stereotype-wise routing, along with effective learning objectives.
  • The first module aims to capture the representation of stereotypes in textual information. Since stereotypes are ubiquitously encoded in recommendation prompts, spanning over word embeddings, it is necessary to effectively distinguish different stereotype groups in addressing stereotype-aware fairness. Accordingly, we introduce a multi-stereotype prompting module to elicit each stereotype group via textual prompts, as recent studies have revealed that LLM stereotypes can be modified by prompts [26, 31, 32].
  • In the second module, we propose a novel stereotype-wise routing strategy over a set of stereotype-relevant experts, to learn unbiased representations. The insight behind lies in first encoding different stereotypes into corresponding experts, then learning adaptive expert weights. Intuitively, the weight distribution is “calibrated” with the original distribution of stereotype groups in user-item interactions.
  • In the last module, learning objectives are carefully designed to facilitate recommendation performance and stereotype-aware fairness of LLM-RS. Notebly, we introduce a stereotype diversity regulation term, to enhance the learning of stereotype information via multiple expert networks.
VI-B Multi-stereotype Prompting
As illustrated in Figure 4, a multi-stereotype prompting module is applied to the textual input prompt of recommendation tasks. The idea of multi-stereotype prompting is to amplify the LLM-encoded stereotypes in recommendation tasks, such as gender, age, and race. In particular, we design stereotype-biased prompts by inserting textual descriptions with respect to each stereotype group, as recent studies have revealed that stereotypes in LLMs can be modified by prompts [26, 31, 32]. For example, a prompting template for gender stereotypes is demonstrated as follows:
[Instruction] Given a female user’s interaction history, identify whether this female user will like the item.
Formally, given a recommendation task prompt $x^{\mathrm{rec}}$ and a set of stereotype-biased prompts $\{x_{i}^{\mathrm{stereotype}}\}_{i=1}^{|\mathcal{G}|}$ (e.g., “female”), the tokenization output of multi-stereotype prompting is formalized as
$x^{\mathrm{rec}},\{x_{i}^{\mathrm{stereotype}}\}_{i=1}^{|\mathcal{G}|}\rightarrow\{(c^{\mathrm{rec}},c_{i}^{\mathrm{stereotype}})\}_{i=1}^{|\mathcal{G}|},$ (6)
@@ VI-B Multi-stereotype Prompting
VI-C Stereotype-wise Routing
@@ VI-C1 Router
$p_{n}(\bm{c}_{i})=\begin{cases}[softmax(\{p_{k}(\bm{c}_{i})\}_{k\in\mathcal{K}})]_{n},&\text{if}\ n=k\\ 0,&\text{otherwise}\end{cases},$ (8)
where $k\in\mathcal{K}$ denotes the index of each activated expert in the top-K set $\mathcal{K}$ (i.e., $|\mathcal{K}|\leq N$). It is worth noting that the routing probabilities of inactivated experts are set to zero, since a reweighting strategy of the routing probabilities will be designed, as illustrated in Section VI-C2. Intuitively, the zeroing operation reinforces the polarization of routing different stereotype information to different experts, facilitating the goal of stereotype-wise routing.
VI-C2 Reweighting Layer
In light of our preliminary findings that the discrimination between user-side and item-side stereotypes can be alleviated in a counterfactual world, we explore combining multiple stereotypes (i.e., stereotype-relevant experts) to address stereotype-aware fairness in LLM-RS. In particular, we aim to generate unbiased representations that are calibrated with the original distribution of stereotypes in user-item interactions, by learning adaptive weights across different stereotype-relevant experts. As a natural solution, the learning objectives of adaptive weights can be obtained by stereotype-aware fairness, as illustrated in Eq. (5), and added to the training loss of LLM-RS. However, unlike discriminative recommendation models, it is challenging to update generative LLM-RS with a learning objective given by group-level fairness [33, 34, 35]. To be specific, the learning objectives of generative LLM-RS are to maximize the likelihood of the label tokens of each output (e.g., target item), which intrinsically fall short in calculating an auxiliary loss over a group of outputs.
@@ VI-C2 Reweighting Layer
$\{w_{n}\}_{n=1}^{N}=\bm{R}(\frac{1}{|\mathcal{G}|}\sum_{i=1}^{|\mathcal{G}|}{\{p_{n}(\bm{c}_{i})\}_{n=1}^{N}}).$ (9)
In other words, $\{w_{n}\}_{n=1}^{N}$ implies a weighted average of multiple stereotypes encoded in each corresponding stereotype-relevant expert.
VI-C3 Prompt-based Experts
With the aforementioned reweighting strategy, the mixture of stereotype-relevant experts can generate unbiased representations against stereotypes, taking advantage of the weighted average of different stereotypes that accord with the original distribution of stereotypes in user-item interactions. To adapt the learned unbiased representations to LLM-RS, we design prompt-based experts to encode the generated representations as soft prompts, which can be seamlessly integrated with LLM-RS via prompt tuning [36]. Formally, the soft prompts generated by experts $\{\bm{E}_{n}\}_{n=1}^{N}$ can be formalized as:
$e=\sum_{n=1}^{N}{w_{n}\bm{E}_{n}(c^{\mathrm{rec}})}.$ (10)
@@ VI-D Learning Objectives
$\mathcal{L}_{\mathrm{rec}}=-\log\Pr_{\Phi+\Theta}(\hat{y}|e,c^{\mathrm{rec}}),$ (12)
@@
$\mathcal{L}_{\mathrm{fair}}\coloneq\mathop{\min}_{\Theta}\|\bm{SF}\|.$ (13)
Notably, we further introduce a stereotype diversity regulation term to enhance the learning of stereotype information via multiple expert networks. In detail, expert parameter redundancy is invertible in multiple-expert architectures, leading to similar representations (i.e., learnable knowledge) across multiple expert networks [37, 38, 39]. Therefore, a stereotype diversity regulation term is designed to maximize the distance between the weights of each stereotype-relevant expert (i.e., diversified representations) as follows:
$\mathop{\max}_{\{w_{n}\}_{n=1}^{N}}\{\mathcal{L}_{\mathrm{expert}}\coloneq\mathop{\min}_{i\neq j}(\|w_{i}-w_{j}\|^{2})\}.$ (14)
@@
$\mathcal{L}_{\mathrm{total}}=\mathcal{L}_{\mathrm{rec}}+\mathcal{L}_{\mathrm{fair}}+\mathcal{L}_{\mathrm{expert}}.$ (15)
VII EXPERIMENT
@@
VII-A Experiment Setup
VII-A1 Datasets
@@
MovieLens100K [40] is a movie recommendation dataset, which provides user-movie interactions and textual information including movie titles and user profiles. In particular, we utilize the binary gender feature of users to assess the gender stereotype of LLM-RS.
BookCrossing [41] is a book recommendation dataset, which provides user-book interactions and textual information including book titles and user profiles. In particular, we utilize the age feature and divide users into teen and adult groups (i.e., under/beyond 18), to assess the age stereotype of LLM-RS.
To maintain a manageable dataset size for efficient LLM-RS training, similar to recent studies [27, 29], we process the original datasets by randomly sampling 10,000 sequences (i.e., each contains 11 chronological interactions). To construct sequential recommendation scenarios, we adopt the leave-one-out strategy and retain the first 10 items in each sequence as the historical interaction, and the last item as the target item. For both datasets, we split the data points of user-item interactions into training, validation, and testing sets with a ratio of 8:1:1, which prevents data leakage. The detailed statistics of experimental datasets are provided in Table III.
TABLE III: Basic statistics of experimental datasets.
@@
  • ::*:∗ : The pair $G_{1}$/$G_{2}$ denotes the male/female stereotype group pair and the adult/teen stereotype group pair, respectively, in MovieLens and BookCrossing.
VII-A2 Evaluation Metrics
We assess the proposed stereotype-aware fairness and the recommendation performance.
Fairness. The proposed stereotype-aware fairness $\bm{SF}$ can be calculated according to Eq. (5). A smaller value of $\bm{SF}$ suggests a more calibrated (i.e., fair) proportion of stereotype groups between recommendation results and the original distribution in user-item interactions.
TABLE IV: Results of fairness-oriented methods. We report the average results over three independent runs with random seeds. The reported results are multiplied by 100, where boldface and underline indicate the best and second best score, respectively. The positive and negative improvements (%) of our proposed method are compared to the best baseline.